OpenStudy (anonymous):

What is the equation in standard form of the line that passes through (-2,6) and has the slope of -1?

OpenStudy (anonymous):

\(\bbox[5pt,skyblue ,border:5px solid blue]{i~can~help~:)}\)

OpenStudy (anonymous):

y=-1x+6 so can you make that into standard now or...

OpenStudy (anonymous):

Not really, i dont understand :| @chrisk123

OpenStudy (anonymous):

ok so let me do it hold on

OpenStudy (anonymous):

(-2,6) you already have the slope, which is -1, so you plug that into the equation which would be: y=-1+b and then you would plug the points in (-2,6) --> 6 = -1(-2) +b 6=2+b -2 -2 b=4 ... so your equation would be y=-1x+4 in standard form that would be x+y=4 thats your answer^^^

OpenStudy (anonymous):

Thank you so much!! (: @chrisk123